Solving for Two Variables using Elimination (must show work)

-4x - 3y = -14 and -x + 3y = -11

Answer by ikleyn(52894)

.

Add the two equations (both sides).


Doing this way, you eliminate "y" and get


    -5x = - 25,  or   x = 5.


Then from the second equation


    3y = -11 + x = -11 + 5 = -6.

     y                     = -6/3 = -2.

ANSWER.  x= 5;  y= -2.

Solved.